OpenStudy (anonymous):

What values of x would make the following expression undefined? x^2-16/x^2-4x-12

myininaya (myininaya):

In other words solve this equation: \[x^2-4x-12=0\]

myininaya (myininaya):

@Somjit please don't just given answers. This is against the Code of Conduct (CoC). http://openstudy.com/code-of-conduct Thank you sir!

myininaya (myininaya):

Can you factor \[x^2-4x-12\] ?

myininaya (myininaya):

oops and I meant give not given

OpenStudy (anonymous):

is that (x-6)(x+2)?

myininaya (myininaya):

Yep yep :)

myininaya (myininaya):

So you have x-6=0 or x+2=0 Solve both and you are done! :)

OpenStudy (anonymous):

but i got x=6 and x=-2. what would you do after then?

myininaya (myininaya):

Yep yep! :)

OpenStudy (anonymous):

thats it?

myininaya (myininaya):

Yep :)

myininaya (myininaya):

The fraction is undefined when the bottom is 0 (remember we don't divide by 0) The bottom is zero when x=6 or x=-2

OpenStudy (anonymous):

k thanks!